A battleship simultaneously fires two shells toward two identical enemy ships. One shell hits ship A, which is close by, and the other hits ship B, which is farther away. The two shells are fired at the same speed. Assume that air resistance is negligible and that the magnitude of the acceleration due to gravity is g.

* What shape is the trajectory (graph of y vs. x) of the shells?
a.straight line
b.parabola
c.hyperbola
d. The shape cannot be determined.

*For two shells fired at the same speed which statement about the horizontal distance traveled is correct?

a. The shell fired at a larger angle with respect to the horizontal lands farther away.
b. The shell fired at an angle closest to 45 degrees lands farther away.
c. The shell fired at a smaller angle with respect to the horizontal lands farther away.
d. The lighter shell lands farther away.

Now, consider for the remaining parts of the question below that both shells are fired at an angle greater than 45 degrees with respect to the horizontal. Remember that enemy ship A is closer than enemy ship B.

* Which shell is fired at the larger angle?

a. A
b. B
c. Both shells are fired at the same angle.

* Which shell is launched with a greater vertical velocity, vy?


a. A
b. B
c. Both shells are launched with the same vertical velocity.


*Which shell is launched with a greater horizontal velocity, vx?

a A
b.B
c.Both shells are launched with the same horizontal velocity.


*Which shell reaches the greater maximum height?

a.A
b.B
c.Both shells reach the same maximum height.

*Which shell has the longest travel time (time elapsed between being fired and hitting the enemy ship)?


a. A
b. B
c. Both shells have the same travel time.
Answer #1
Concepts and reason

The concept used to solve the problem is body undergoing projectile motion.

First, find the relation between the position coordinates by using the projectile motion equations.

The horizontal distance travelled maximum height of the projectile and travel time can be computed using projectile equations.

Fundamentals

The shells fired from a battleship undergo projectile motion. The shape of trajectory for a projectile motion is always parabola as the object fired at some angle from the horizontal travels with same horizontal velocity as air resistance is negligible and vertically the gravity pulls it down with acceleration due to gravitygg.

The equation of trajectory for a projectile is,

rf=ri+vit+12gt2{\vec r_f} = {\vec r_i} + {\vec v_i}t + \frac{1}{2}\vec g{t^2}

Here, ri{\vec r_i} is the initial position vector, vi{\vec v_i} is the initial velocity vector, g\vec g is the acceleration due to gravity, and tt is any arbitrary time.

The xxcomponent and yycomponent of the position vector of the projectile are,

x=vitcosθy=vitsinθ12gt2\begin{array}{c}\\x = {v_i}t\cos \theta \\\\y = {v_i}t\sin \theta - \frac{1}{2}g{t^2}\\\end{array}

Here, vi{v_i} is the initial velocity, gg is the acceleration due to gravity, tt is any arbitrary time, and θ\theta is the angle which initial velocity make with respect to horizontal.

The general equation of parabola is,

y=ax2y = a{x^2}

Here, xx and yy are the coordinates, and aandba{\rm{ and }}b are any constants such that a>0a > 0.

The general equation of a hyperbola is,

x2a2y2b2=1\frac{{{x^2}}}{{{a^2}}} - \frac{{{y^2}}}{{{b^2}}} = 1

Here, xx and yy are the coordinates, and aandba{\rm{ and }}b are any constants such that a>0,b>0andaba > 0,b > 0{\rm{ and }}a \ne b.

The general equation of straight line is,

y=mx+cy = mx + c

Here, xx and yy are the coordinates, mm is the slope, and cc is the intercept.

The xandyx{\rm{ and }}y components of the initial velocity is,

vxi=vicosθvyi=visinθ\begin{array}{l}\\{v_{xi}} = {v_i}\cos \theta \\\\{v_{yi}} = {v_i}\sin \theta \\\end{array}

Here, vxi{v_{xi}} is the xx-component of the initial velocity, vyi{v_{yi}} is the yy-component of the initial velocity, vi{v_i} is the initial speed, and θ\theta is the angle which initial velocity make with respect to horizontal.

The time of flight or time for which projectile moves is,

T=2visinθgT = \frac{{2{v_i}\sin \theta }}{g}

The maximum height for a projectile is,

h=vi2sin2θ2gh = \frac{{v_i^2{{\sin }^2}\theta }}{{2g}}

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The maximum value of sine function is 1 at an angle 9090^\circ . The value of cosine is more for small angles and sine is more for large angles.

(a)

The equation of trajectory for a projectile is,

rf=ri+vit+12gt2{\vec r_f} = {\vec r_i} + {\vec v_i}t + \frac{1}{2}\vec g{t^2}

Here, ri{\vec r_i} is the initial position vector, vi{\vec v_i} is the initial velocity vector, g\vec g is the acceleration due to gravity, and tt is any arbitrary time.

Use the xx component of the position coordinate and solve for time.

x=vitcosθt=xvicosθ\begin{array}{c}\\x = {v_i}t\cos \theta \\\\t = \frac{x}{{{v_i}\cos \theta }}\\\end{array}

Substitute t=xvicosθt = \frac{x}{{{v_i}\cos \theta }} in the equation y=vitsinθ12gt2y = {v_i}t\sin \theta - \frac{1}{2}g{t^2} and find the relation between xx and yy coordinates.

y=vi(xvicosθ)sinθ12g(xvicosθ)2=xsinθcosθ12g(x2vi2cos2θ)\begin{array}{c}\\y = {v_i}\left( {\frac{x}{{{v_i}\cos \theta }}} \right)\sin \theta - \frac{1}{2}g{\left( {\frac{x}{{{v_i}\cos \theta }}} \right)^2}\\\\ = x\frac{{\sin \theta }}{{\cos \theta }} - \frac{1}{2}g\left( {\frac{{{x^2}}}{{v_i^2{{\cos }^2}\theta }}} \right)\\\end{array}

Compare the equation with equation of circle it is clear that the trajectory is not a straight line.

The equation of trajectory for a projectile is,

rf=ri+vit+12gt2{\vec r_f} = {\vec r_i} + {\vec v_i}t + \frac{1}{2}\vec g{t^2}

Here, ri{\vec r_i} is the initial position vector, vi{\vec v_i} is the initial velocity vector, g\vec g is the acceleration due to gravity, and tt is any arbitrary time.

Use the xx component of the position coordinate and solve for time.

x=vitcosθt=xvicosθ\begin{array}{c}\\x = {v_i}t\cos \theta \\\\t = \frac{x}{{{v_i}\cos \theta }}\\\end{array}

Substitute t=xvicosθt = \frac{x}{{{v_i}\cos \theta }} in the equation y=vitsinθ12gt2y = {v_i}t\sin \theta - \frac{1}{2}g{t^2} and find the relation between xx and yy coordinates.

y=vi(xvicosθ)sinθ12g(xvicosθ)2=xsinθcosθ12g(x2vi2cos2θ)\begin{array}{c}\\y = {v_i}\left( {\frac{x}{{{v_i}\cos \theta }}} \right)\sin \theta - \frac{1}{2}g{\left( {\frac{x}{{{v_i}\cos \theta }}} \right)^2}\\\\ = x\frac{{\sin \theta }}{{\cos \theta }} - \frac{1}{2}g\left( {\frac{{{x^2}}}{{v_i^2{{\cos }^2}\theta }}} \right)\\\end{array}

Compare the equation with equation of parabola it is clear that the trajectory is a parabola.

The equation of trajectory for a projectile is,

rf=ri+vit+12gt2{\vec r_f} = {\vec r_i} + {\vec v_i}t + \frac{1}{2}\vec g{t^2}

Here, ri{\vec r_i} is the initial position vector, vi{\vec v_i} is the initial velocity vector, g\vec g is the acceleration due to gravity, and tt is any arbitrary time.

Use the xx component of the position coordinate and solve for time.

x=vitcosθt=xvicosθ\begin{array}{c}\\x = {v_i}t\cos \theta \\\\t = \frac{x}{{{v_i}\cos \theta }}\\\end{array}

Substitute t=xvicosθt = \frac{x}{{{v_i}\cos \theta }} in the equation y=vitsinθ12gt2y = {v_i}t\sin \theta - \frac{1}{2}g{t^2} and find the relation between xx and yy coordinates.

y=vi(xvicosθ)sinθ12g(xvicosθ)2=xsinθcosθ12g(x2vi2cos2θ)\begin{array}{c}\\y = {v_i}\left( {\frac{x}{{{v_i}\cos \theta }}} \right)\sin \theta - \frac{1}{2}g{\left( {\frac{x}{{{v_i}\cos \theta }}} \right)^2}\\\\ = x\frac{{\sin \theta }}{{\cos \theta }} - \frac{1}{2}g\left( {\frac{{{x^2}}}{{v_i^2{{\cos }^2}\theta }}} \right)\\\end{array}

Compare the equation with equation of parabola it is clear that the trajectory is not a hyperbola.

The equation of trajectory for a projectile is,

rf=ri+vit+12gt2{\vec r_f} = {\vec r_i} + {\vec v_i}t + \frac{1}{2}\vec g{t^2}

Here, ri{\vec r_i} is the initial position vector, vi{\vec v_i} is the initial velocity vector, g\vec g is the acceleration due to gravity, and tt is any arbitrary time.

Use the xx component of the position coordinate and solve for time.

x=vitcosθt=xvicosθ\begin{array}{c}\\x = {v_i}t\cos \theta \\\\t = \frac{x}{{{v_i}\cos \theta }}\\\end{array}

Substitute t=xvicosθt = \frac{x}{{{v_i}\cos \theta }} in the equation y=vitsinθ12gt2y = {v_i}t\sin \theta - \frac{1}{2}g{t^2} and find the relation between xx and yy coordinates.

y=vi(xvicosθ)sinθ12g(xvicosθ)2=xsinθcosθ12g(x2vi2cos2θ)\begin{array}{c}\\y = {v_i}\left( {\frac{x}{{{v_i}\cos \theta }}} \right)\sin \theta - \frac{1}{2}g{\left( {\frac{x}{{{v_i}\cos \theta }}} \right)^2}\\\\ = x\frac{{\sin \theta }}{{\cos \theta }} - \frac{1}{2}g\left( {\frac{{{x^2}}}{{v_i^2{{\cos }^2}\theta }}} \right)\\\end{array}

Thus, the trajectory can be determined.

(b)

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The maximum value of sine function is 1 at an angle 9090^\circ . So, the horizontal distance is maximum when the angle is 9090^\circ .

2θ=90θ=45\begin{array}{c}\\2\theta = 90^\circ \\\\\theta = 45^\circ \\\end{array}

The shell fired at angles larger than 4545^\circ will travel less distance as compared to the shells fired at 4545^\circ . As the value of sine function will decrease after and before 4545^\circ . The horizontal distance travelled is dependent on sine function only as shells are fired at same speed and gg is same. Therefore, shells fired at a larger angle with respect to the horizontal lands closer to the battleship and not farther away.

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The maximum value of sine function is 1 at an angle 9090^\circ . So, the horizontal distance is maximum when the angle is 9090^\circ .

2θ=90θ=45\begin{array}{c}\\2\theta = 90^\circ \\\\\theta = 45^\circ \\\end{array}

The shell fired at angles closer to 4545^\circ will travel more distance as compared to the shells fired at angles far more or less than 4545^\circ . As the value of sine function will decrease after and before 4545^\circ . The horizontal distance travelled is dependent on sine function only as shells are fired at same speed and gg is same. Therefore, shells fired at an angle closest to 4545^\circ lands farther away.

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The maximum value of sine function is 1 at an angle 9090^\circ . So, the horizontal distance is maximum when the angle is 9090^\circ .

2θ=90θ=45\begin{array}{c}\\2\theta = 90^\circ \\\\\theta = 45^\circ \\\end{array}

The shell fired at angles smaller than 4545^\circ will travel less distance as compared to the shells fired at 4545^\circ . As the value of sine function will decrease after and before 4545^\circ . The horizontal distance travelled is dependent on sine function only as shells are fired at same speed and gg is same. Therefore, shells fired at a smaller angle with respect to the horizontal lands closer to the battleship and not farther away.

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The horizontal distance does not depend on the mass of the shells. The shells of different mass but with same speed and fired at same angle will land at same distance. Therefore, the lighter shell does not land farther away.

(c)

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The enemy ship A is closer than enemy ship B. This means the shell A has travelled smaller horizontal distance. The shell fired at angles larger than 4545^\circ will travel less distance as compared to the shells fired at 4545^\circ . As the value of sine function will decrease as angles goes more and more large after 4545^\circ . The horizontal distance travelled is dependent on sine function only as shells are fired at same speed and gg is same. Therefore, shells fired at a larger angle with respect to the horizontal lands closer to the battleship. Hence, the shell A is fired at larger angle.

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The enemy ship A is closer than enemy ship B. This means the shell A has travelled smaller horizontal distance. The shell fired at angles larger than 4545^\circ will travel less distance as compared to the shells fired at 4545^\circ . As the value of sine function will decrease as angles goes more and more large after 4545^\circ . The horizontal distance travelled is dependent on sine function only as shells are fired at same speed and gg is same. Therefore, shells fired at a larger angle with respect to the horizontal lands closer to the battleship. Thus, the shell B is not fired at larger angle.

The horizontal distance travelled by a projectile is,

R=vi2sin2θgR = \frac{{v_i^2\sin 2\theta }}{g}

The enemy ship A is closer than enemy ship B. This means the shell A has travelled smaller horizontal distance. The shell fired at angles larger than 4545^\circ will travel less distance as compared to the shells fired at 4545^\circ . As the value of sine function will decrease as angles goes more and more large after 4545^\circ . The horizontal distance travelled is dependent on sine function only as shells are fired at same speed and gg is same. Therefore, shells fired at a larger angle with respect to the horizontal lands closer to the battleship. Thus, the shells are fired at different angle.

(d)

The yy or vertical component of the initial velocity is,

vyi=visinθ{v_{yi}} = {v_i}\sin \theta

Here, vyi{v_{yi}} is the yy-component of the initial velocity, vi{v_i} is the initial speed, θ\theta is the angle which initial velocity make with respect to horizontal.

The vertical component of the velocity depends on sine of the angle launched. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell A is launched at larger angle. Hence, the vertical velocity vy{v_y} is greater of the A shell.

The yy or vertical component of the initial velocity is,

vyi=visinθ{v_{yi}} = {v_i}\sin \theta

Here, vyi{v_{yi}} is the yy-component of the initial velocity, vi{v_i} is the initial speed, θ\theta is the angle which initial velocity make with respect to horizontal.

The vertical component of the velocity depends on sine of the angle launched. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell B is launched at smaller angle. Hence, the vertical velocity vy{v_y} is smaller of the B shell.

The yy or vertical component of the initial velocity is,

vyi=visinθ{v_{yi}} = {v_i}\sin \theta

Here, vyi{v_{yi}} is the yy-component of the initial velocity, vi{v_i} is the initial speed, θ\theta is the angle which initial velocity make with respect to horizontal.

The vertical component of the velocity depends on sine of the angle launched. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell A and B are launched at different angles. Hence, the vertical velocity vy{v_y} is different for both shells.

(e)

The xx or horizontal component of the initial velocity is,

vxi=vicosθ{v_{xi}} = {v_i}\cos \theta

Here, vxi{v_{xi}} is the xx-component of the initial velocity, vi{v_i} is the initial speed, θ\theta is the angle which initial velocity make with respect to horizontal.

The horizontal component of the velocity depends on cosine of the angle launched. The cosine is smaller for larger angles till it reaches minimum at 9090^\circ . The shell A is launched at larger angle. Hence, the horizontal velocity vx{v_x} is smaller for the A shell.

The xx or horizontal component of the initial velocity is,

vxi=vicosθ{v_{xi}} = {v_i}\cos \theta

Here, vxi{v_{xi}} is the xx-component of the initial velocity, vi{v_i} is the initial speed, θ\theta is the angle which initial velocity make with respect to horizontal.

The horizontal component of the velocity depends on cosine of the angle launched. The cosine is smaller for larger angles till it reaches maximum at 9090^\circ . The shell A is launched at larger angle. Hence, the horizontal velocity vx{v_x} is greater for the B shell.

The xx or horizontal component of the initial velocity is,

vxi=vicosθ{v_{xi}} = {v_i}\cos \theta

Here, vxi{v_{xi}} is the xx-component of the initial velocity, vi{v_i} is the initial speed, θ\theta is the angle which initial velocity make with respect to horizontal.

The horizontal component of the velocity depends on cosine of the angle launched. The cosine is smaller for larger angles till it reaches maximum at 9090^\circ . The shell A and B are launched at different angles. Hence, the horizontal velocity vx{v_x} is different for both shells.

(f)

The maximum height for a projectile is,

h=vi2sin2θ2gh = \frac{{v_i^2{{\sin }^2}\theta }}{{2g}}

The maximum height depends on square of the sine of the angle launched only as shells are fired at same speed and gg is same. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell A is launched at larger angle. Hence, the maximum height is greater of the A shell.

The maximum height for a projectile is,

h=vi2sin2θ2gh = \frac{{v_i^2{{\sin }^2}\theta }}{{2g}}

The maximum height depends on square of the sine of the angle launched only as shells are fired at same speed and gg is same. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell B is launched at smaller angle. Hence, the maximum height is smaller of the B shell.

The maximum height for a projectile is,

h=vi2sin2θ2gh = \frac{{v_i^2{{\sin }^2}\theta }}{{2g}}

The maximum height depends on square of the sine of the angle launched only as shells are fired at same speed and gg is same. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell A and B are launched at different angles. Hence, the maximum height is different for both shells.

(g)

The time of flight or travel time for which projectile moves is,

T=2visinθgT = \frac{{2{v_i}\sin \theta }}{g}

The travel time of the shell depends on sine of the angle launched only as both shells are launched at same speed and gg is also same. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell A is launched at larger angle. Hence, the travel time is greater for the shell A.

The time of flight or travel time for which projectile moves is,

T=2visinθgT = \frac{{2{v_i}\sin \theta }}{g}

The travel time of the shell depends on sine of the angle launched only as both shells are launched at same speed and gg is also same. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell B is launched at smaller angle. Hence, the travel time is smaller for the shell B.

The time of flight or travel time for which projectile moves is,

T=2visinθgT = \frac{{2{v_i}\sin \theta }}{g}

The travel time of the shell depends on sine of the angle launched only as both shells are launched at same speed and gg is also same. The sine is more for large angles till it reaches maximum at 9090^\circ . The shell A and B are launched at different angles. Hence, the travel time is different for both shells.

Ans: Part a

The shape of the trajectory is parabola.